OpenStudy (anonymous):

If f(x) = |x – 5| + 3, find f(2).

OpenStudy (anonymous):

I got 6.

jimthompson5910 (jim_thompson5910):

f(x) = |x – 5| + 3 f(2) = |2 – 5| + 3 f(2) = |-3| + 3 f(2) = 3 + 3 f(2) = 6 so you are correct. Nice job

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Thank you. :)

OpenStudy (mathmale):

I agree: you're right. Do consider sharing the work you do to get from one point to another, as it's easier to give you meaningful feedback if I can observe your work towards a solution.

OpenStudy (anonymous):

I did exactly what jim did. I'll post my work on the next question.

OpenStudy (mathmale):

Great. Happy to see you building up momentum and understanding what to do!